OpenStudy (anonymous):

Use the five steps for problem solving to answer the following question. Please show all of your work. Enrollment at a particular online university is 11,600. There are seven times as many students who work full time as there are students who do not. How many students work full time and how many do not?

OpenStudy (anonymous):

You need to break this down into a mathematical equation. Lets look at the first line Enrollment at a particular online university is 11,600, which is the total students enrolled = 11,600 Lets break down the seconded sentence, which is below There are seven times as many students who work full time as there are students who do not Lets look at the first part, which is below There are seven times as many students who work full time = 7x Lets look at the second part, which is below as there are students who do not work full time = x Now lets look at the last sentence How many students work full time and how many do not? Ok lets find how many students do not work full time 7x + x = 11,600 8x = 11,600 8x/8 = 11,600/8 x = 1450 do not work full time Now lets find out how many work full time. Remember , there are seven times as many students who work full time f(x) = 7x f(1450) = 7x or 1450 * 7 = 10150 people work full time